OpenStudy (anonymous):

Fun Land theme park has a total of 140 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? 40 60 100 110

OpenStudy (johnweldon1993):

Set up the equation... 140 total so males + females = 140 the number of males is 20 more than twice of what the females are so M = 2F + 20 M + F = 140 we know M = 2F + 20 so 2F + 20 + F = 140 can you solve for F?

OpenStudy (johnweldon1993):

No problem we'll walk through it....do you understand what I have done so far??

OpenStudy (anonymous):

yessss

OpenStudy (johnweldon1993):

Okay great...then you have a basic understanding of how to do it! so we're at 2F + 20 + F = 140 We want to solve for F (females) after that we solve for M (Males like the question asks 2F + 20 + F = 140 we combine like terms 2F + 1F = 3F right? so we have 3F + 20 = 140 now lets subtract 20 from both sides...this will isolate 3F 3F + 20 = 140 -20 -20 3F = 120 now divide both sides by 3 to solve for F 3F = 120 --------- 3 3 F = 40 right Make sense so far?

OpenStudy (anonymous):

ohhhhhhhhhhhhhh okay thank you so much (:

OpenStudy (johnweldon1993):

Not quite done yet...that solved for Females...the question asked for males....so how many males are there?

OpenStudy (anonymous):

120?

OpenStudy (johnweldon1993):

If there are 140 people in total...and 40 are female...how many are male?

OpenStudy (anonymous):

100

OpenStudy (johnweldon1993):

There you go...did that make sense or do you need more explanation?

OpenStudy (anonymous):

it makes sense (: